OpenStudy (anonymous):

At Cheng's Bike Rentals, it cost $22 to rent a bike for 5 hours.How many hours of bike use does a customer get per dollar?

OpenStudy (phi):

they give you a ratio of \[\frac{$22}{5 \text{ hours} }\] they want to find the ratio \[\frac{\text{ hours}} {$1}\] so the first thing to do is flip the ratio they gave you, so you have hours per dollars \[\frac{5 \text{ hours} }{$22}\] now divide top and bottom by 22 to make the denominator 1 what do you get for the top?

OpenStudy (anonymous):

0.227?

OpenStudy (phi):

yes, you get 5/22 or about 0.227 hours of course nobody talks about 0.227 hours (though that is the answer here): 0.227 hours per dollar is the answer but in minutes, 60*0.227 = 13.6 minutes 13 and 1/2 minutes per dollar sounds more "normal"